# Glued lattices are better quantizers than $K_{12}$

### v2

@@ Abstract
40 years ago, Conway and Sloane proposed using the highly symmetrical Coxeter–Todd lattice $K_{12}$ for quantization, and estimated its second moment. Since then, all published lists identify $K_{12}$ as the best 12-dimensional lattice quantizer. Surprisingly, $K_{12}$ is not optimal: we construct two new 12-dimensional lattices with lower normalized second moments. The new lattices are obtained by gluing together 6-dimensional lattices.
@@ I Introduction
The above-mentioned tables devote more attention to some dimensions than others. One such dimension is $n=12$, and the reason for the interest in this dimension is the existence of the highly symmetrical Coxeter–Todd lattice $K_{12}$, which was discovered by Coxeter and Todd in 1953 [17]. Its second moment was computed numerically in 1984 [12] and it is listed as the best known 12-dimensional lattice quantizer in all tables we have seen. The lattice was conjectured optimal for both quantization and packing in [1, p. 13]. Its construction, symmetry group, and other properties are described in [18], [1, pp. 127–129].
@@ IV-A Glued $E_{6}\times E_{6}$
confirming the numerical estimate in Table I. Its covariance matrix is proportional to the identity, which is a necessary but not sufficient condition for global and local optimality [28, 16]. The complete face catalog is available online as supplementary data [29].
@@ IV-B Glued $D_{6}\times D_{6}$
This glue group is the point group $C_{2v}$, which has order four. The group identity is ${\boldsymbol{g}}_{0}$, and the multiplication table reads ${\boldsymbol{g}}_{1}^{2}={\boldsymbol{g}}_{2}^{2}={\boldsymbol{g}}_{3}^{2}={\boldsymbol{g}}_{0}$, ${\boldsymbol{g}}_{1}{\boldsymbol{g}}_{2}={\boldsymbol{g}}_{3}$, ${\boldsymbol{g}}_{1}{\boldsymbol{g}}_{3}={\boldsymbol{g}}_{2}$, and ${\boldsymbol{g}}_{2}{\boldsymbol{g}}_{3}={\boldsymbol{g}}_{1}$, where the (Abelian) group operation is addition modulo $D_{6}$.
@@
We consider all subsets $\Gamma$ of the glue words $(D_{6}^{*}/D_{6})\times(D_{6}^{*}/D_{6})$ such that $\tilde{L}$ in (6) is a lattice, i.e., subsets that are groups under addition modulo $L$. There are $67$ such subsets. However, several of these subsets are equivalent to each other, in the sense that a rotation and/or reflection operation in the symmetry group of $L$ transforms all elements of one subset into the element of another subset. The relevant symmetry operations are (i) interchanging ${\boldsymbol{g}}_{ij}$ with ${\boldsymbol{g}}_{ji}$ throughout $\Gamma$, i.e. swapping the first set of $6$ coordinates with the last $6$; (ii) replacing all occurrences of ${\boldsymbol{g}}_{1j}$ with ${\boldsymbol{g}}_{3j}$ and vice versa; and (iii) replacing all occurrences of ${\boldsymbol{g}}_{i1}$ with ${\boldsymbol{g}}_{i3}$ and vice versa. If only inequivalent subsets of glue words are considered, $22$ subsets remain. We furthermore exclude the $10$ subsets that are direct products of $6$-dimensional glue groups, which generate product lattices. The remaining $12$ glue groups $\Gamma$ are listed in Table II. The NSM of each corresponding lattice $\tilde{L}$ was estimated as in Section IV-A.
@@
Using the same method as before, it was found that the Voronoi region has $1912$ facets, 21 273 4562127345621\,273\,45621 273 456 vertices, and 10 395 549 5531039554955310\,395\,549\,55310 395 549 553 faces overall. The faces fall into 2 54225422\,5422 542 equivalence classes under its symmetry group, which has order 1 061 683 20010616832001\,061\,683\,2001 061 683 200. The face catalog is available online as supplementary data [29]. The NSM of the new lattice is
@@ V Conclusions
As a byproduct, an improved $13$-dimensional lattice quantizer is obtained. The previously best reported lattice for $n=13$ is a product of $K_{12}$ and a scaled integer lattice $a{\mathbb{Z}}$, with an NSM of $0.071035$ [15, Tab. I]. Replacing $K_{12}$ with the new best $12$-dimensional lattice in a similar product construction yields a slightly improved NSM of $0.070974$.

### v3

@@ Abstract
40 years ago, Conway and Sloane proposed using the highly symmetrical Coxeter–Todd lattice $K_{12}$ for quantization, and estimated its second moment. Since then, all published lists identify $K_{12}$ as the best 12-dimensional lattice quantizer. Surprisingly, $K_{12}$ is not optimal: we construct two new 12-dimensional lattices with lower normalized second moments. The new lattices are obtained by gluing together products of two 6-dimensional lattices.
@@ I Introduction
The above-mentioned tables devote more attention to some dimensions than others. One such dimension is $n=12$, and the reason for the interest in this dimension is the existence of the highly symmetrical Coxeter–Todd lattice $K_{12}$, which was discovered by Coxeter and Todd in 1953 [17]. Its second moment was computed numerically in 1984 [12] and exactly in 2009 [13]. It is listed as the best $12$-dimensional lattice quantizer in all tables we have seen. The lattice was conjectured optimal for both quantization and packing in [1, p. 13]. Its construction, symmetry group, and other properties are described in [18], [1, pp. 127–129].
@@ III Gluing Theory
More generally, the construction (6) is also valid when $\Gamma$ is not a subset of $L^{*}/L$. For example, the $9$-dimensional lattice quantizer with the smallest known NSM is $L\cup(L+{\boldsymbol{g}})$ (corresponding to $\Gamma=\{{\boldsymbol{0}},{\boldsymbol{g}}\}$), where $L=D_{8}\times 2a{\mathbb{Z}}$, ${\boldsymbol{g}}$ is a deep hole of $L$, and $a$ is an algebraic scalar constant [11, 6]. In this case, $L$ is not an integer lattice and $L^{*}/L$ does not exist. While interesting, such generalizations of (6) are not considered further in this paper.
@@ IV-A Glued $E_{6}\times E_{6}$
confirming the numerical estimate in Table I. Its covariance matrix is proportional to the identity, which is a necessary but not sufficient condition for global and local optimality [28, 16]. The complete face catalog is available online [29, Ancillary files].
@@ IV-B Glued $D_{6}\times D_{6}$
This glue group is the point group $C_{\mathrm{2v}}$, which has order four. The group identity is ${\boldsymbol{g}}_{0}$, and the multiplication table reads ${\boldsymbol{g}}_{1}^{2}={\boldsymbol{g}}_{2}^{2}={\boldsymbol{g}}_{3}^{2}={\boldsymbol{g}}_{0}$, ${\boldsymbol{g}}_{1}{\boldsymbol{g}}_{2}={\boldsymbol{g}}_{3}$, ${\boldsymbol{g}}_{1}{\boldsymbol{g}}_{3}={\boldsymbol{g}}_{2}$, and ${\boldsymbol{g}}_{2}{\boldsymbol{g}}_{3}={\boldsymbol{g}}_{1}$, where the (Abelian) group operation is addition modulo $D_{6}$.
@@
We consider all subsets $\Gamma$ of the glue words $(D_{6}^{*}/D_{6})\times(D_{6}^{*}/D_{6})$ such that $\tilde{L}$ in (6) is a lattice, i.e., subsets that are groups under addition modulo $L$. There are $67$ such subsets. However, several of these subsets are equivalent to each other, in the sense that a rotation and/or reflection operation in the symmetry group of $L$ transforms all elements of one subset into the elements of another subset. The relevant symmetry operations are (i) interchanging ${\boldsymbol{g}}_{ij}$ with ${\boldsymbol{g}}_{ji}$ throughout $\Gamma$, i.e., swapping the first set of $6$ coordinates with the last $6$; (ii) replacing all occurrences of ${\boldsymbol{g}}_{1j}$ with ${\boldsymbol{g}}_{3j}$ and vice versa; and (iii) replacing all occurrences of ${\boldsymbol{g}}_{i1}$ with ${\boldsymbol{g}}_{i3}$ and vice versa. If only inequivalent subsets of glue words are considered, $22$ subsets remain. We furthermore exclude the $10$ subsets that are direct products of $6$-dimensional glue groups, which generate product lattices. The remaining $12$ glue groups $\Gamma$ are listed in Table II. The NSM of each corresponding lattice $\tilde{L}$ was estimated as in Section IV-A.
@@
Using the same method as before, it was found that the Voronoi region has $1912$ facets, 21 273 4562127345621\,273\,45621 273 456 vertices, and 10 395 549 5531039554955310\,395\,549\,55310 395 549 553 faces overall. The faces fall into 2 54225422\,5422 542 equivalence classes under its symmetry group, which has order 1 061 683 20010616832001\,061\,683\,2001 061 683 200. The face catalog is available online [29, Ancillary files]. The NSM of the new lattice is
@@ V Conclusions
As a byproduct, an improved $13$-dimensional lattice quantizer is obtained. The previously best published lattice for $n=13$ is a product of $K_{12}$ and a scaled integer lattice $a{\mathbb{Z}}$, with an NSM of $0.071035$ [15, Tab. I]. Replacing $K_{12}$ with the new best $12$-dimensional lattice in a similar product construction yields a slightly improved NSM of $0.070974$ (which is however inferior to a yet unpublished laminated lattice [27]).